(y+z)p+(z+x)q=x+y


1
Expert's answer
2022-02-10T14:29:07-0500

"\displaystyle
(y+z)p+(z+x)q=x+y"

The auxiliary equations are;

"\displaystyle
\frac{dx}{y+z}=\frac{dy}{z+x}=\frac{dz}{x+y}"

From the first two equation, we have;

"\displaystyle
\frac{dx}{y+z}=\frac{dy}{z+x}\\"

integrating both sides yields;

"\displaystyle
\Rightarrow\frac{x}{y+z}=\frac{y}{z+x}+a", where a is an arbitrary constant


"\displaystyle
\Rightarrow\frac{x}{y+z}-\frac{y}{z+x}=a\cdots\cdots\cdots\cdots(1)"

From the last two equation, we have;

"\displaystyle
\frac{dy}{z+x}=\frac{dz}{x+y}"

integrating both sides;

"\displaystyle
\frac{y}{z+x}=\frac{z}{x+y}+b", where a is an arbitrary constant


"\displaystyle
\Rightarrow\frac{y}{z+x}-\frac{z}{x+y}=b\cdots\cdots\cdots\cdots(2)"


from (1) and (2), we have the solution as;

"\displaystyle
f\left(\frac{x}{y+z}-\frac{y}{z+x},\ \frac{y}{z+x}-\frac{z}{x+y}\right)=0"